# Enhancing Robustness of Vision-Language Models through Orthogonality Learning and Cross-Regularization

### v1

@@ 4.2 Comparison with other methods
More importantly, our method surpasses the comparative $LoRA_{CLIP}$ with 2.74%, 6.15% and 4.95% of base, novel and HM evaluation, which further demonstrates the OrthCR is capable of not only efficiently adapting to task-specific task but also leading to generalizability preservation, thanks to the norm-preserving property of orthogonal finetuning. And these results further presents the prevalent $LoRA_{CLIP}$ method potentially tends to prioritize task-specific knowledge and results in task overfitting issues while ours has no such issues, especially for the few-shot image recognition task. Meanwhile, our approach reports consistent superorities beyond the conventional prompt learning methods, VPT and IVLP, better illustrate the effectiveness of our approach. When compared with competing MaPLe [37] and PromptSRC [38] which utilize complex strategies to enhance prompt tuning, our method still behaves better generalizability, obtaining highest accuracy on evaluation with 76.55% for new classes and 80.02% for HM.
Cross-dataset transfer. For evaluating the cross-dataset tranfer, we train our approach on ImageNet [67] and then directly evaluate it on the other datasets without any domain-specific finetuning or adaptation. We compare cross-dataset performance with existing methods in Table 4. In comparison with CoOp [88] and CoCoOp [87], our proposed OrthCR presents better generalization performance in 9/10 and 5/10 datasets, respectively. Importantly, our approach exceeds $LoRA_{CLIP}$ in 9/10 datasets and shows obvious advantages among these dataset, which further demonstrates that our methods retains stronger zero-shot generalizability. Meanwhile, compared with the prompt tuning methods MaPLe [37] and PromptSRC [38], we obtain 7/10 and 6/10 better generalization performance while not introducing any tunable parameters after training (0 v.s. 3.55MB and 0 v.s 46KB, respectively) and no complicated training strategy tailored to struggle with the generalizability preservation.
Domain generalization. Table 3 reports the results of OrthCR and other methods on out-of-distribution datasets. Following the common methods, we train our model and directly evaluate on other datasets. We can observe that our method consistently surpasses $LoRA_{CLIP}$ on all datasets, while obtaining 3/4 superiority with CoOp and CoCoOp. Interestingly, prompt-based VPT illustrates inferior performance in 4/4 datasets to ours, while ours gains 2/4 better generlization evaluation beyond MaPLe [37]. This suggests that our orthogonal tuning with simple yet effective cross-regularization enables the finetuned model favor better generalization for datasets with domain shifts.
@@ 4.3 Ablations and analysis
Methods CoOp CoCoOp VPT PLOT MAPLE OrthCR
@@ 5 Conclusions
This paper proposes a novel and efficient method for adapting pretrained VLM weights, OrthCR, for specific downstream tasks (e.g., few-shot image recognition). To explore an effective finetuning approach not suffering from task overfitting issues under a data-efficient setting, we propose an orthogonal finetuning method for efficiently updating pretrained weights. Optimized by the constraint with Cayley parameterization during training, the finetuned CLIP model is capable of maintaining minimal and same-level of hyperspherical energy as the pretrained model owing to norm-preserving property, leading to better robustness and generalizability for task-specific scenarios. Meanwhile, a cross-regularization strategy is designed to enforce the model not to deviate far away from the pretrained one within a bypass manner. Additionally, we first exploit the Cutout data augmentation to enable the finetuned model to learn better task-specific knowledge on a small data set. Finally, extensive experiments demonstrate the training efficiency and generalizability preservation of our approach and showcase competitive performance on three generalization evaluations, shedding new light on the future works for this few-shot tuning task.
@@ A.1 More implementation details
Input Interpolation "Bicubic"
@@
Transforms ["random resized crop", "random filp", "normalize"]
@@
Warmup Type "constant"
@@
Prompt Initialization "a photo of a"
Precision "fp16"
@@ A.2 Evaluation metrics
Among all our experiments, we report $top_{1}$ accuracy for each dataset. In base-to-base/base-to-new generalization, the $top_{1}$ accuracy is measured on base classes and new classes, respectively. We then calculate the harmonic mean (HM) between the base and new class accuracy to show the generalization trade-off [78], using $HM=\frac{2\times base\times new}{base+new}$. In domain generalization, and cross-dataset transfer settings, we measure $top-1$ accuracy on the test set of each dataset with the same split provided by CoOp [88] following other related works.
@@ A.3 More dataset descriptions
We throughly conduct our method on publicly available 15 image recognition datasets across 4 common generalizability evaluation settings: ImageNet [67] and Caltech101 [19] for generic objects classification, Oxford_Pets [60], StanfordCars [39], Flowers102 [58], Food101 [5] and FGVCAircraft [55] for fine-grained classification, SUN397 [79] for scene recognition, DTD [14] for texture classification, EuroSAT [27] for satellite imagery recognition and UCF101 [71] for action recognition; datasets with apparent domain shifts ImageNetV2 [66], ImageNet-Sketch [75], ImageNet-A [30] and ImageNet-R [29]. We make a summary in terms of data statistics in Table 7.

### v3

@@ 4.2 Comparison with other methods
More importantly, our method surpasses the comparative $LoRA_{CLIP}$ with 2.74%, 6.15% and 4.95% of base, novel and HM evaluation, which further demonstrates the OrthSR is capable of not only efficiently adapting to task-specific task but also leading to generalizability preservation, thanks to the norm-preserving property of orthogonal finetuning. And these results further presents the prevalent $LoRA$ method potentially tends to prioritize task-specific knowledge and results in task overfitting issues while ours has no such issues, especially for the few-shot image recognition task. Meanwhile, our approach reports consistent superorities beyond the conventional prompt learning methods, VPT and IVLP, better illustrate the effectiveness of our approach. When compared with competing MaPLe Khattak et al. (2023a) and PromptSRC Khattak et al. (2023b) which utilize complex strategies to enhance prompt tuning, our method still behaves better generalizability, obtaining highest accuracy on evaluation with 76.55% for new classes and 80.02% for HM.
Cross-dataset transfer. For evaluating the cross-dataset tranfer, we train our approach on ImageNet Russakovsky et al. (2015) and then directly evaluate it on the other datasets without any domain-specific finetuning or adaptation. We compare cross-dataset performance with existing methods in Table 4. In comparison with CoOp Zhou et al. (2022b) and CoCoOp Zhou et al. (2022a), our proposed OrthSR presents better generalization performance in 9/10 and 5/10 datasets, respectively. Importantly, our approach exceeds $LoRA_{CLIP}$ in 9/10 datasets and shows obvious advantages among these dataset, which further demonstrates that our methods retains stronger zero-shot generalizability. Meanwhile, compared with the prompt tuning methods MaPLe Khattak et al. (2023a) and PromptSRC Khattak et al. (2023b), we obtain 7/10 and 6/10 better generalization performance while not introducing any tunable parameters after training (0 v.s. 3.55MB and 0 v.s 46KB, respectively) and no complicated training strategy tailored to struggle with the generalizability preservation.
Domain generalization. Table 3 reports the results of OrthSR and other methods on out-of-distribution datasets. Following the common methods, we train our model and directly evaluate on other datasets. We can observe that our method consistently surpasses $LoRA_{CLIP}$ on all datasets, while obtaining 3/4 superiority with CoOp and CoCoOp. Interestingly, prompt-based VPT illustrates inferior performance in 4/4 datasets to ours, while ours gains 2/4 better generlization evaluation beyond MaPLe Khattak et al. (2023a). This suggests that our orthogonal tuning with simple yet effective cross-regularization enables the finetuned model favor better generalization for datasets with domain shifts.
@@ 4.3 Ablations and analysis
Methods CoOp CoCoOp VPT PLOT MAPLE OrthSR
@@ 5 Conclusions
This paper proposes a novel and efficient method for adapting pretrained VLM weights, OrthSR, for specific downstream tasks (e.g., few-shot image recognition). To explore an effective fine-tuning approach not suffering from task overfitting issues under a data-efficient setting, we propose an orthogonal fine-tuning method for efficiently updating pretrained weights. Optimized by the constraint with Cayley parameterization during training, the fine-tuned CLIP model is capable of maintaining minimal and same-level of hyperspherical energy as the pretrained model owing to norm-preserving property, leading to better robustness and generalizability for task-specific scenarios. Meanwhile, a self-regularization strategy is designed to enforce the model not to deviate far away from the pretrained one within a bypass manner. Additionally, we first explore attentive CutOut data augmentation to enable the fine-tuned model to learn better task-specific knowledge on a small data set. Finally, extensive experiments demonstrate the training efficiency and generalizability preservation of our approach and showcase competitive performance on three generalization evaluations, shedding new light on the future works for this few-shot tuning task.
@@ A.1 More implementation details
Input Interpolation ”Bicubic”
@@
Transforms [”random resized crop”, ”random filp”, ”normalize”]
@@
Warmup Type ”constant”
@@
Prompt Initialization ”a photo of a”
Precision ”fp16”
@@ A.2 Evaluation metrics
Among all our experiments, we report $top_{1}$ accuracy for each dataset. In base-to-base/base-to-new generalization, the $top_{1}$ accuracy is measured on base classes and new classes, respectively. We then calculate the harmonic mean (HM) between the base and new class accuracy to show the generalization trade-off Xian et al. (2017), using $HM=\frac{2\times base\times new}{base+new}$. In domain generalization, and cross-dataset transfer settings, we measure $top-1$ accuracy on the test set of each dataset with the same split provided by CoOp Zhou et al. (2022b) following other related works.
@@ A.3 More dataset descriptions
We throughly conduct our method on publicly available 15 image recognition datasets across 4 common generalizability evaluation settings: ImageNet Russakovsky et al. (2015) and Caltech101 Fei-Fei (2004) for generic objects classification, Oxford_Pets Parkhi et al. (2012), StanfordCars Krause et al. (2013), Flowers102 Nilsback & Zisserman (2008), Food101 Bossard et al. (2014) and FGVCAircraft Maji et al. (2013) for fine-grained classification, SUN397 Xiao et al. (2010) for scene recognition, DTD Cimpoi et al. (2014) for texture classification, EuroSAT Helber et al. (2019) for satellite imagery recognition and UCF101 Soomro et al. (2012) for action recognition; datasets with apparent domain shifts ImageNetV2 Recht et al. (2019), ImageNet-Sketch Wang et al. (2019), ImageNet-A Hendrycks et al. (2021b) and ImageNet-R Hendrycks et al. (2021a). We make a summary in terms of data statistics in Table 7.
@@ A.4 Loss balancing hyper-parameters sensitivity ablations
Appendix B Theoretical Proof
Following previous works Chen et al. (2024); Ni et al. (2024), this section provides detailed proofs for the Theorem in Sec. 3.6. Notably, we propose to utilize attentive CutOut data augmentation to implicitly increase the sample number and make use of pre-trained model as generalization  anchor to maintain the generalization error bound, which is different from  Chen et al. (2024). We introduce the following lemmas for proving our Theorem.
Lemma 1(McDiarmid’s Inequality Vershynin (2018)). Consider independent random variables $v_{1},v_{2},\cdots,v_{n}\in\mathcal{V}$ and a function $\phi:\mathcal{V}^{n}\to\mathbb{R}$. Suppose that for all $v_{1},v_{2},\cdots,v_{n}$ and ${v_{i}}^{\prime}\in\mathcal{V}\left(i=1,2,\cdots,n\right)$, the function satisfies
$\left|\phi\left(v_{1},\cdots,V_{i-1},V_{i},V_{i+1},\cdots,V_{n}\right)-\phi\left(v_{1},\cdots,V_{i-1},{v_{i}}^{\prime},V_{i+1},\cdots,V_{n}\right)\right|\leq c_{i},\vspace{0pt}$ (13)
and then it holds that
$\mathcal{P}\left\{\phi\left(v_{1},v_{2},\cdots,v_{n}\right)-\mathbb{E}_{v_{1},v_{2},\cdots,v_{n}}\left(\phi\left(v_{1},v_{2},\cdots,v_{n}\right)\right)>\mu\right\}\leq e^{-\frac{2\mu^{2}}{{\textstyle\sum_{i=1}^{n}c_{i}^{2}}}}.\vspace{0pt}$ (14)
The proof of Theorem 1. is given as follows.
Theorem 1. Assume that $\Theta^{*}$ is the solution to OrthSR. Then we have that for any $10<\varepsilon<$ 1 with probability $1-\varepsilon$,
$\epsilon(\Theta^{*})-\bar{\epsilon}_{\chi}(\Theta^{*})\leq X^{*}\sqrt{\frac{2\ln(1/\delta)}{N}}+\frac{C^{\prime\prime}}{\lambda^{2\alpha}\sqrt{N}}.$
where $\epsilon(\Theta^{*})$ is the true error. $\bar{\epsilon}_{\chi}(\Theta^{*})$ is the empirical error. $X^{*}$ is the upper bound of the loss function $L$. $N$ is the number of training samples. $\lambda$ is our introduced regularization parameter. $\alpha>0$. $\delta$ is a probability parameter. $C^{\prime\prime}$ encompasses constants from the Rademacher complexity bound.
Proof.
The generalization error is defined as:
$\epsilon(\Theta)=\mathbb{E}_{(x,y)\sim D}\left[L(s_{\Theta}(x),y)\right]$
where $\Theta$ represents the model parameters, $L(s_{\Theta}(x),y)$ is the loss function, and $D$ is the true data distribution.
The empirical error is:
$\bar{\epsilon}_{\chi}(\Theta)=\frac{1}{N}\sum_{i=1}^{N}L(s_{\Theta}(x_{i}),y_{i})$
where $\chi=\{(x_{i},y_{i})\}_{i=1}^{N}$ is the training set, and $N$ is the sample size.
We use McDiarmid’s inequality to control the deviation between empirical error and true error. The inequality states:
$P\left(f(X_{1},\ldots,X_{n})-\mathbb{E}[f(X_{1},\ldots,X_{n})]>t\right)\leq\exp\left(-\frac{2t^{2}}{\sum_{i=1}^{n}c_{i}^{2}}\right)$
where $X_{1},X_{2},\ldots,X_{n}$ are independent random variables, and $f(X_{1},\ldots,X_{n})$ is a function of these variables. When one sample in the training set changes, the maximum change in the empirical error is:
$\Delta=\bar{\epsilon}_{\chi}(\Theta)-\bar{\epsilon}_{\chi^{\prime}}(\Theta)$
The change in empirical error is bounded by $\frac{c}{N}$, where $c$ is the upper bound on the difference in the loss function:
$|L(s_{\Theta}(x),y)-L(s_{\Theta}(x^{\prime}),y^{\prime})|\leq c$
Applying McDiarmid’s inequality with the bound $\frac{c}{N}$, we obtain the following bound:
$P\left(\epsilon(\Theta)-\bar{\epsilon}_{\chi}(\Theta)>t\right)\leq\exp\left(-\frac{2Nt^{2}}{c^{2}}\right)$
We introduce the Rademacher complexity $R_{N}(L)$, which measures the complexity of the model:
$R_{N}(L)=\mathbb{E}_{\sigma,\chi}\left[\sup_{\Theta\in\mathcal{H}}\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}L(s_{\Theta}(x_{i}),y_{i})\right]$
The generalization error bound becomes:
$\epsilon(\Theta)\leq\bar{\epsilon}_{\chi}(\Theta)+2R_{N}(L)+X^{*}\sqrt{\frac{2\ln(1/\delta)}{N}}$
where: $\bar{\epsilon}_{\chi}(\Theta)$ is the empirical error. $2R_{N}(L)$ is the Rademacher complexity term. $X^{*}\sqrt{\frac{2\ln(1/\delta)}{N}}$ is the variance term that decreases as the sample size $N$ increases. To further reduce the generalization error, we introduce the regularization term $L_{KD}$ (Knowledge Distillation Loss) in Eq. 10, which limits the complexity of the model. The objective function of our OrthSR is:
$\min_{\Theta}\left(L_{CE}+\lambda L_{KD}\right)$
where $L_{CE}$ is the cross-entropy loss for measuring the fit of the model. $L_{KD}$ is the knowledge distillation loss, reducing the difference between student and teacher models. $\lambda$ controls the trade-off between the two losses. To understand why the Rademacher complexity $R_{N}(L)$ is reduced under the regularization term, we analyze how regularization influences the hypothesis space $\mathcal{H}$ and, consequently, the complexity of the loss function class.
The Rademacher complexity $R_{N}(L)$ measures the richness of the loss class =$\mathcal{L}=\{L(s_{\Theta}(x),y):\Theta\in\mathcal{H}\}$ by evaluating how well it can fit random noise. It is defined as:
$R_{N}(L)=\mathbb{E}_{\sigma,\chi}\left[\sup_{\Theta\in\mathcal{H}}\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}L(s_{\Theta}(x_{i}),y_{i})\right],$
where $\sigma_{i}$ are independent Rademacher variables taking values ±$\pm 1$ with equal probability.
Regularization introduces a penalty term $\lambda L_{KD}$ in the objective function:
$\min_{\Theta}\left(L_{CE}+\lambda L_{KD}\right).$
This penalty discourages complex models by imposing a cost on large parameter values or deviations from the teacher model in knowledge distillation. As a result, the effective hypothesis space $\mathcal{H}_{\lambda}$ becomes smaller or more restricted because models with high complexity are penalized.
Mathematically, stronger regularization (larger $\lambda$) enforces tighter constraints on $\Theta$, effectively reducing the norm or other measures of complexity of the model parameters. We assume that through regularization, the model parameters satisfy the following constraint:
$\|\Theta\|\leq\frac{C}{\lambda^{\beta}},$
where $C$ and $\beta>0$ are constants.
Under this constraint, and assuming that the loss function $L$ is Lipschitz continuous with Lipschitz constant $L_{0}$, the Rademacher complexity can be bounded as:
$R_{N}(L)\leq\frac{L_{0}C^{\prime}}{\lambda^{\beta}\sqrt{N}},$
where $C^{\prime}$ is another constant.
Substituting this bound into the generalization error bound, we have:
$\epsilon(\Theta^{*})-\bar{\epsilon}_{\chi}(\Theta^{*})\leq X^{*}\sqrt{\frac{2\ln(1/\delta)}{N}}+\frac{1}{\lambda^{\alpha}}\cdot R_{N}(L)\leq X^{*}\sqrt{\frac{2\ln(1/\delta)}{N}}+\frac{L_{0}C^{\prime}}{\lambda^{\alpha+\beta}\sqrt{N}}.$
To ensure consistency in the exponents of $\lambda$, we set:
$\alpha=\beta>0.$
Therefore, the generalization error bound becomes:
$\epsilon(\Theta^{*})-\bar{\epsilon}_{\chi}(\Theta^{*})\leq X^{*}\sqrt{\frac{2\ln(1/\delta)}{N}}+\frac{C^{\prime\prime}}{\lambda^{2\alpha}\sqrt{N}},$
where $C^{\prime\prime}=L_{0}C^{\prime}$ is a constant.
This inequality shows that $R_{N}(L)$ decreases as $\lambda$ increases, since $\alpha>0$. By reducing $R_{N}(L)$ through regularization, we tighten the generalization error bound:
$\epsilon(\Theta^{*})-\bar{\epsilon}_{\chi}(\Theta^{*})\leq X^{*}\sqrt{\frac{2\ln(1/\delta)}{N}}+\frac{C^{\prime\prime}}{\lambda^{2\alpha}\sqrt{N}}.$
In summary, the regularization term reduces the Rademacher complexity $R_{N}(L)$ by limiting the capacity of the hypothesis space $\mathcal{H}$. This reduction leads to better generalization performance by preventing overfitting and tightening the generalization error bound.